OpenStudy (anonymous):

9/a=3/8 solve each proportion.

OpenStudy (anonymous):

cross multiply to get 9*8 = 3*a 72 = 3*a 72/3 = a 24 = a

OpenStudy (anonymous):

cross multiply 3a = 9 x 8 a = 72/3 = 24